Problem 17.5

After emptying her lungs, a person inhales 4.0 L of air at 2.0 ∘C and holds her breath.

Part A

How much does the volume of the air increase as it warms to her body temperature of 36 ∘C ?

Express your answer using two significant figures.
Answer #1

Given , V1 = 4 L , T1 = 2°C = 275 K , T2 = 36°C = 309°C

Since the pressure is constant ,

V1/T1 = V2/T2

4/275 = V2/309

V2 = 4*309/275

V2 = 4.495 L

Thus Increase in the volume = V2 - V1 = (4.495 - 4)L

= 0.495 L